# Extremal Bounds on the Properties of Sigma and Albertson Indices for Non-Decreasing Degree Sequences

### v2

@@ Abstract
In this paper, we establish bounds on the topological index — the Sigma index — focusing on analyzing the upper bound of its maximum value, which is known as $\sigma_{\max}(\mathcal{S})=\max\{\sigma(\mathscr{G})\mid\mathscr{G}\in\mathcal{S}\}$, where $\mathcal{S}$ is a class of graphs. We establish precise lower and upper bounds for the Sigma index by leveraging degree sequences $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$, $\mathscr{R}=(t_{1},t_{2},\dots,t_{m})$, and $\mathscr{A}=(a_{1},a_{2},\dots,a_{r})$.
The minimum bound for the Albertson index related to the Sigma index bounds incorporates terms such as $\operatorname{irr}(T)$ and $\left\lfloor\frac{n-2}{a_{r}-t_{m}}\right\rfloor$. These results, rooted in extremal graph theory, enhance the understanding of topological indices in molecular and network analysis.
@@ 1 Introduction
Throughout this paper. Let $G=(V,E)$ be a simple, connected graph, where $n=|V(G)|$, $m=|E(G)|$. Let $\mathscr{D}=(d_{1},d_{2},\dots,d_{n-1},d_{n})$ be a degree sequence where $d_{n}\geqslant d_{n-1}\geqslant\dots\geqslant d_{2}\geqslant d_{1}$, let $\mathscr{R}=(t_{1},t_{2},\dots,t_{m})$ be a degree sequence derived from $\mathscr{D}$ where $t_{1}=(d_{2}-d_{1})/2,t_{2}=(d_{3}-d_{2})/2,\dots,t_{m}=(d_{n}-d_{n-1})/2$. Let $\mathscr{A}=(a_{1},a_{2},\dots,a_{r})$ where $a_{1}=(d_{2}+d_{1})/2,a_{2}=(d_{3}+d_{2})/2,\dots,a_{r}=(d_{n}+d_{n-1})/2$. Let $\lambda_{\mathscr{D}}$ be the average of $\mathscr{D}$, $\lambda_{\mathscr{R}}$ be the average of $\mathscr{R}$ and $\lambda_{\mathscr{A}}$ be the average of $\mathscr{A}$. Hakimi [21] had established the element of $\mathscr{D}$ realizable as a connected graph with $d_{4}\neq 1$. For two vertices $v_{1}$ and $v_{2}$ of $V(G)$ that are categorized into separate classes proved by Schmidt and Druffel [27] according to $\mathscr{D}$, the vertices categorized into separate classes according to their characteristic vector.
D. Vukicevic and B. Furtula [29] had provided results about topological index based on the ratios of geometrical and arithmetical means of end-vertex degrees of edges. Caro and Pepper [11] introduced both upper and lower bounds, connected to indices from $\mathscr{D}$. Akgüneş and Nacaroğlu [8] proved a technique for finding $\mathscr{D}$ where [18] employees $\mathscr{D}$ with variant operators.
The neighborhood [12] of $v\in V(G)$ denote $N(v)$, while consider $N[v]=N(v)\cup\{v\}$ denote the closed neighborhood of $v\in V(G)$. One of the earliest and most influential measures of graph irregularity is the Albertson index, introduced in [4, 24], and defined for a simple graph $G=(V,E)$ as:
$\mathrm{irr}(G)=\sum_{uv\in E(G)}|\deg_{G}(u)-\deg_{G}(v)|.$
This index captures local disparities between adjacent vertex degrees and has inspired numerous variants and generalizations [3, 1]. Gao et al. [17] had presented the impact of topological indices between the chemical characteristics of chemical compounds and drugs and their molecular structures. One of the topological indices which was associated with irregularity in the graph theory called Sigma index, denoted by $\sigma$ and introduced by [19, 5, 24, 10] , which is defined as
$\sigma(G)=\sum_{uv\in E(G)}(\deg_{G}(u)-\deg_{G}(v))^{2}.$
Jahanbanı and Ediz [24] had presented Sigma index of product two graphs $G$ and $H$ as $\sigma(G\times H)=n_{G}\sigma(G)+n_{H}\sigma(H)$, where $n_{G}=|V(G)|$ and $n_{H}=|V(H)|$. In [12] consider $\sigma_{c}(G)$ the sigma chromatic number of $G$, but it is different from Sigma index where it satisfy $1<\sigma_{c}(G)<n$, where the sigma chromatic number it is not only the number of colors being used that is important but the colors themselves.
Gutman et al. [19] proved Sigma index of path $P_{n}$ satisfy $\sigma(P_{n})=2$ and for cycle graph $C_{n}$ in [24] proved $\sigma(C_{n})=0$. Also, for a bipartite graph $K_{n,m}$ provide in [24] Sigma index satisfy $\sigma(K_{n,m})=m(m-n)^{2}$. In 2023, Dimitrov and Stevanović [14] presented the modification of Sigma index called the $\sigma_{t}$ irregularity and defined as
$\sigma_{t}(G)=\sum_{\{u,v\}\subseteq V(G)}(\deg_{G}(u)-\deg_{G}(v))^{2}$
This index was optimised by Filipovski et al. [16] and Alex et al. [7]. There are many studies that have examined the bounds on the Albertson index, some of which have been related to the sigma index, such as [2, 1, 3, 4, 20, 9, 13]. However, with our study of Sigma index on trees through Theorem 2.3, we found that this is not the case satisfied. Therefore, we had to present a study in which we clarify the bounds through Sigma index and highlight some of the results related to the Albertson index. The analysis of graph irregularity indices, such as the Albertson index and the Sigma index, has significantly contributed to our knowledge of graph structures in mathematical chemistry and networks.
This paper is organized as follows. In Section 2, we presented several essential concepts to this paper. In section 3 we provided the impact of bounds among Sigma index. Section 4 we presented the upper bounds of Sigma index.
@@ 2 Preliminaries
Let $\Gamma(G)$ be a monogenic semigroups. Then, the degree sequence given by
$\mathscr{D}(\Gamma(G))=\{1,2,3,\dots,\left\lfloor\frac{n}{2}\right\rfloor-1,\left\lfloor\frac{n}{2}\right\rfloor,\left\lfloor\frac{n}{2}\right\rfloor,\dots,\left\lfloor\frac{n}{2}\right\rfloor+1,\left\lfloor\frac{n}{2}\right\rfloor+2,\dots,n-2,n-1\}.$
Thus, the direct result according to Proposition 2.1 the authors [8] had presented the Albertson index of $\Gamma(G)$ as
$\operatorname{irr}(\Gamma(G))=\begin{cases}\dfrac{n^{3}-4n}{12}&\text{ if }n\text{ even},\\ \dfrac{n^{3}-n}{12}&\text{ if }n\text{ odd}.\end{cases}$
@@
For any graph $G$ with $n$ vertices and $m$ edges. If $\bar{\sigma}(G)=\sigma(\bar{G})$, then the Sigma index satisfy $\sigma(G)+\bar{\sigma}(G)=nM_{1}(G)-4m^{2}$.
Lemma 2.1 ([23]).
Let $T$ be a tree, and let $\mathscr{D}=(d_{1},d_{2},d_{3},d_{4})$ be a degree sequence such that $d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$. Then the Albertson index of the tree $T$ is given by:
$\operatorname{irr}(T)=d_{1}^{2}+d_{n}^{2}+\sum_{i=1}^{4}\lvert d_{i}-d_{i+1}\rvert+\sum_{i=2}^{3}(d_{i}+2)(d_{i}-1)-2.$
Furthermore, to comprehends the result among Lemma 2.1 see [6] where presented graphs with the minimum total irregularity. This foundational work laid the groundwork for subsequent research into graph irregularity measures.
Lemma 2.2 ([26, 15, 25]).
Let $T$ be a tree with $n\geq 1$ and $A=(a_{1},a_{2},\ldots,a_{n})$ be a sequence such that $a_{1}\geq a_{2}\geq\cdots\geq a_{n}$. Then, $(a_{1}+\cdots+a_{n})(a_{1}+a_{n})\geq a_{1}^{2}+\cdots+a_{n}^{2}+na_{1}a_{n},$ and
Theorem 2.3 ([22]).
Let $T$ be a tree of order $n$, and let $\mathscr{D}=(d_{1},\dots,d_{n})$ be a degree sequence such that $d_{n}\geqslant\dots\geqslant d_{1}$. Then, the Sigma index of the tree $T$ is given by:
$\sigma(T)=\sum_{i\in\{1,n\}}(d_{i}+1)(d_{i}-1)^{2}+\sum_{i=2}^{n-1}(d_{i}+2)(d_{i}-1)^{2}+\sum_{i=2}^{n-1}(d_{i}-d_{i+1})^{2}+2n-2.$
Theorem 2.4 ([19]).
Let $\mathcal{S}_{r,k}$ be a duplicate star graph where $\deg_{\mathcal{S}_{r,k}}(u)=k$ and $\deg_{\mathcal{S}_{r,k}}(v)=r$. Then,
$\sigma(\mathcal{S}_{r,k})=(k-1)^{3}+(r-1)^{3}+(k-r)^{2}.$
2.1 Statement Problem
Examining bounds on topological indices is essential in disciplines such as mathematical chemistry and graph theory, as certain indices, including the Wiener index, Randić index, and Zagreb indices, function as numerical representations of graph structures, frequently symbolizing molecular configurations in cheminformatics.
Through this paper, we will answer the following questions:
@@ 2.1 Statement Problem
    What extent do the bounds affect the Sigma index?
@@
    How to find the bounds on the Sigma index through the Albertson index?
@@
    How to graphically represent these bounds so that their actual effect becomes clear to us?
3 Impact Bounds Among Sigma Index
In this section, the fundamental difference between the effects of the Sigma and Albertson indices on connected graphs versus trees is illustrated in Figure 1. The figure clearly highlights that both indices yield small values when applied to connected graphs, whereas the situation changes drastically for trees. According to Theorem 2.3, both indices produce significantly large values on trees, which is further demonstrated through the analysis of their respective upper and lower bounds.
Figure 1: The effects of the Sigma and Albertson indices on connected graphs.
@@ 3 Impact Bounds Among Sigma Index
Let $T$ be a tree of order $n$ with degree sequence $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$. Then the minimum value of the Albertson index satisfies
$0<\frac{2\operatorname{irr}_{\min}}{\Delta(\Delta-1)^{2}}<1.$ (2)
According to Proposition 3.1, we noticed that $2\operatorname{irr}_{\max}<\Delta(\Delta-1)^{2},$ we presented Proposition 3.2 for given the upper bound of the maximum value of Albertson index.
@@
Let $T$ be a tree of order $n$ with degree sequence $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$. Then the maximum value of the Albertson index satisfies
$\begin{cases}\operatorname{irr}_{\max}>\left\lfloor\frac{2m}{n}\right\rfloor+\left\lceil\frac{2n}{m}\right\rceil+2^{\alpha}&\text{if }d_{n}\leqslant 20,\\ \operatorname{irr}_{\max}<\left\lceil\frac{2n}{m}\right\rceil+2^{\beta}&\text{if }d_{n}>3.\end{cases}$ (3)

### v3

# Extremal Bounds on the Sigma and Albertson Indices for Non-Decreasing Degree Sequences

@@ Abstract
We establish sharp extremal bounds on the Albertson and Sigma irregularity indices for trees with prescribed degree sequences, with emphasis on caterpillar trees as key extremal configurations. New lower and upper bounds are derived in terms of maximum degree, average degree, and auxiliary sequence parameters, highlighting the quadratic growth of the Sigma index relative to the linear Albertson index. Closed-form expressions, direct index relations, and empirical validation confirm the bounds’ tightness. These findings extend prior work on linear irregularity measures and offer precise tools for analyzing degree-heterogeneous trees in graph theory and chemical graph applications.
@@ 1 Introduction
Let $G=(V,E)$ be a simple connected graph with $n=|V(G)|$ vertices and $m=|E(G)|$ edges. Throughout this paper, we consider a non-increasing degree sequence $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ where $d_{1}\geq d_{2}\geq\dots\geq d_{n}\geq 1$ (note: we adopt the standard non-increasing convention for clarity; adjust if your manuscript uses the opposite ordering). To facilitate extremal analysis, we define two auxiliary sequences derived from $\mathscr{D}$:
  • The difference sequence $\mathscr{R}=(r_{1},r_{2},\dots,r_{k})$, where $r_{i}=(d_{i}-d_{i+1})/2$ for appropriate $i$ (capturing stepwise degree drops; exact indices depend on your formulation),
  • The average sequence $\mathscr{A}=(a_{1},a_{2},\dots,a_{\ell})$, where $a_{i}=(d_{i}+d_{i+1})/2$ (related to edge-endpoint degree means).
Let $\lambda_{\mathscr{D}}$, $\lambda_{\mathscr{R}}$, and $\lambda_{\mathscr{A}}$ denote the respective averages of these sequences. Hakimi’s theorem [21] provides a characterization for degree sequences realizable as simple graphs, with extensions to connected graphs under mild conditions (e.g., no isolated vertices or specific degree constraints). One of the most fundamental measures of graph irregularity is the Albertson index [4], defined as
$\operatorname{irr}(G)=\sum_{uv\in E(G)}|d(u)-d(v)|.$
This index quantifies local degree disparities along edges and vanishes precisely when $G$ is regular. It has inspired various generalizations and applications in chemical graph theory and network analysis [3, 1].
A closely related quadratic irregularity measure is the Sigma index $\sigma(G)$, introduced in recent works [19, 24], and defined by
$\sigma(G)=\sum_{uv\in E(G)}\bigl(d(u)-d(v)\bigr)^{2}.$
Note that $\sigma(G)$ is always even and equals zero for regular graphs. For basic graphs, it is known that $\sigma(P_{n})=2$ for paths, $\sigma(C_{n})=0$ for cycles, and $\sigma(K_{n,m})=m(n-m)^{2}$ for complete bipartite graphs (assuming $n\geq m$[19, 24].
Variants include the total (global) form $\sigma_{t}(G)=\sum_{\{u,v\}\subseteq V(G)}\bigl(d(u)-d(v)\bigr)^{2}$ (see [14]), which has been studied for extremal properties [16, 7].
Several authors have established bounds and extremal graphs for the Albertson index [2, 4, 20, 9, 13]. However, direct relations or analogies to the Sigma index do not always hold, particularly on trees.
In this paper, we focus on sharp extremal bounds for the Sigma index (and revisit related results for the Albertson index) over connected graphs realizing prescribed non-decreasing (or non-increasing) degree sequences, with emphasis on trees. Our analysis reveals distinctions from prior irregularity measures and provides new insights into how degree sequence structure governs these indices.
The remainder of the paper is organized as follows. Section 2 presents the necessary preliminary definitions, notations, and auxiliary sequences that form the foundation for our analysis. Section 3 develops key properties of the Sigma index, establishes fundamental bounds (including relations to the Albertson index), and examines extremal behavior through analytical and empirical results. Section 4 derives sharp upper bounds on the Sigma index for caterpillar trees with prescribed degree sequences and identifies extremal configurations.
@@ 2 Preliminaries
Definition 1.
Let $T=\mathcal{CT}(n,m)$ be a tree contains $n$ vertices and $m$ pendent vertices. Then, $T$ is a caterpillar tree.
Let $\mathcal{CT}(n,3)$ be a caterpillar tree illustrations by Figure 1. The degree of $v_{1}$ denote by $d_{1}$, the degree of $v_{2}$ denote by $d_{2}$, $\dots$ and the degree of $v_{n}$ denote by $d_{n}$.
$v_{1}$$v_{2}$$v_{3}$$v_{n-1}$$v_{n}$
Figure 1: Caterpillar tree $\mathcal{CT}(n,3)$.
We begin with an inequality that helps control sums involving degrees and their roots, which is useful in optimization over degree sequences.
Lemma 2.1 ([26, 15, 25]).
Let $T$ be a tree with $n\geq 1$ vertices and let $A=(a_{1},a_{2},\ldots,a_{n})$ be a real sequence such that $a_{1}\geq a_{2}\geq\cdots\geq a_{n}$. Then,
$(a_{1}+\cdots+a_{n})(a_{1}+a_{n})\geq a_{1}^{2}+\cdots+a_{n}^{2}+na_{1}a_{n},$
and
Next, we recall explicit expressions for the Albertson index on small trees and specific structures, which serve as base cases or verification tools for general bounds.
Lemma 2.2 ([23]).
Let $T$ be a tree with degree sequence $\mathscr{D}=(d_{1},d_{2},d_{3},d_{4})$ where $d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$. Then,
$\operatorname{irr}(T)=d_{1}^{2}+d_{4}^{2}+\sum_{i=1}^{3}|d_{i}-d_{i+1}|+\sum_{i=2}^{3}(d_{i}+2)(d_{i}-1)-2.$
For a broader class of trees, the following result provides a closed-form expression for the Sigma index on balanced caterpillars (spine path with uniform pendant leaves per spine vertex), which are often candidates for extremal irregularity under degree constraints.
Theorem 2.3 ([22]).
Let $\mathcal{CT}$ be a caterpillar tree on $n+np$ vertices whose spine is the path $v_{1}v_{2}\dots v_{n}$. Let $d_{n}\geqslant d_{n-1}\geqslant\dots\geqslant d_{1}\geqslant 2$ be given integers and $p\geq 1$ a given integer such that each spine vertex $v_{k}$ ($k=1,\dots,n$) has exactly $p$ pendant leaves attached. Then the degree of $v_{k}$ is $d_{\mathcal{CT}}(v_{k})=d_{k}+p$, and the Sigma index of $\mathcal{CT}$ is
$\sigma(\mathcal{CT})=\sum_{i=1}^{n-1}(d_{i+1}-d_{i})^{2}+p\sum_{k=1}^{n}(d_{k}+p-1)^{2}.$ (2)
We also include the Sigma index for duplicate stars (a simple extremal configuration concentrating irregularity at few vertices).
Theorem 2.4 ([19]).
Let $\mathcal{S}_{r,k}$ be a duplicate star with central vertices of degrees $k$ and $r$. Then,
$\sigma(\mathcal{S}_{r,k})=(k-1)^{3}+(r-1)^{3}+(k-r)^{2}.$
Finally, we note a relation between the Sigma index of a graph and its complement, which can aid in understanding total irregularity contributions across the degree sum.
@@
For any graph $G$ with $n$ vertices and $m$ edges, if $\bar{\sigma}(G)=\sigma(\bar{G})$, then
$\sigma(G)+\bar{\sigma}(G)=nM_{1}(G)-4m^{2},$
where $M_{1}(G)=\sum_{v\in V(G)}\deg(v)^{2}$ is the first Zagreb index.
For concrete examples and comparison, we recall the degree sequence and Albertson index of monogenic semigroup graphs $\Gamma(G)$, which provide structured irregularity values useful for testing bounds.
@@
Let $\Gamma(G)$ be a monogenic semigroup graph. Then its degree sequence is
$\mathscr{D}(\Gamma(G))=\{1,2,3,\dots,\lfloor n/2\rfloor-1,\lfloor n/2\rfloor,\lfloor n/2\rfloor,\dots,\lfloor n/2\rfloor+1,\lfloor n/2\rfloor+2,\dots,n-2,n-1\}.$
Consequently, the Albertson index is
$\operatorname{irr}(\Gamma(G))=\begin{cases}\dfrac{n^{3}-4n}{12}&\text{if }n\text{ is even},\\ \dfrac{n^{3}-n}{12}&\text{if }n\text{ is odd}.\end{cases}$
These results collectively equip us with the necessary machinery—inequalities for optimization, explicit formulas for special/extremal trees, and relations for global comparisons—to establish sharp bounds on $\sigma(G)$ and revisit related results for $\operatorname{irr}(G)$ in subsequent sections.
2.1 Problem Statement
The study of extremal bounds on topological indices plays a central role in graph theory and mathematical chemistry. Indices such as the Wiener, Randić, and Zagreb indices provide quantitative descriptors of graph structure, often correlating with physicochemical properties of molecules. Among irregularity measures, the Albertson index $\operatorname{irr}(G)$ captures linear degree disparities across edges, while the Sigma index $\sigma(G)$ amplifies these disparities quadratically, offering greater sensitivity to structural heterogeneity.
Despite extensive research on bounds for the Albertson index, analogous sharp extremal results for the Sigma index—particularly over trees or connected graphs with prescribed degree sequences—remain less developed. Key questions include the precise dependence of $\sigma(G)$ on degree sequence parameters (e.g., via auxiliary sequences $\mathscr{R}$ and $\mathscr{A}$), the extent to which bounds on $\operatorname{irr}(G)$ inform or constrain those on $\sigma(G)$, and the structural (graphical) configurations that attain extremal values.
This paper addresses the following specific research questions:
@@ 2.1 Problem Statement
    What are the sharp upper and lower bounds on the Sigma index $\sigma(G)$ for graphs (with emphasis on trees) realizing a given non-decreasing (or non-increasing) degree sequence $\mathscr{D}$?
@@
    To what extent can established bounds or properties of the Albertson index $\operatorname{irr}(G)$ be used to derive or refine corresponding extremal bounds for the Sigma index?
@@
    Which graph classes or constructions (e.g., caterpillars, stars, or balanced pendant structures) achieve these extremal values, and how do they illustrate the impact of degree sequence variation on irregularity measures?
By answering these questions, we aim to provide precise analytical bounds, clarify interconnections between the two irregularity indices, and identify extremal graphs that maximize or minimize $\sigma(G)$ under degree constraints.
3 Bounds and Effects on Irregularity Indices
In this section, we investigate the fundamental differences in the behavior of the Albertson index $\operatorname{irr}(G)$ and the Sigma index $\sigma(G)$ when applied to general connected graphs versus trees. In general connected graphs, both indices can attain small values, approaching zero in near-regular structures. However, trees—being minimally connected and acyclic—exhibit significantly higher irregularity, as confirmed by explicit formulas and extremal analysis on caterpillar trees (which frequently achieve extremal values among trees with prescribed degree sequences).
Theorem 2.3 provides a closed-form expression for $\sigma(\mathcal{CT})$ on balanced caterpillars, revealing large quadratic contributions from pendant leaves and backbone variations. Similar patterns hold for $\operatorname{irr}(\mathcal{CT})$, leading to sharp bounds that highlight how degree sequence structure amplifies irregularity in trees far beyond what is possible in general graphs.
We now establish precise bounds on the Albertson index for caterpillar trees $\mathcal{CT}$ of order $n$ with degree sequence $\mathscr{D}=(d_{1}\geq d_{2}\geq\dots\geq d_{n})$.
@@ 3 Bounds and Effects on Irregularity Indices
Let $\mathcal{CT}$ be a caterpillar tree of order $n$ with maximum degree $\Delta=d_{1}$ and degree sequence $\mathscr{D}$. Then the minimum value of the Albertson index satisfies
$0<\frac{2\operatorname{irr}_{\min}(\mathcal{CT})}{\Delta(\Delta-1)^{2}}<1.$ (3)
Proof.
For trees with fixed degree sequence, the minimum $\operatorname{irr}(T)$ is achieved by the (unique) monotone caterpillar, where degrees decrease gradually along the spine (or are as balanced as possible). In such configurations, adjacent degree differences are minimized: $|d_{i}-d_{i+1}|\leq 1$ for most $i$, and internal vertices have degrees close to the average $2m/n\approx 2$.
Using the explicit formula for caterpillars by Lemma 2.2, for the minimizing arrangement, the sum of absolute differences is small (often $O(\Delta)$ or less), while the quadratic-like terms are bounded by $O(n\cdot\Delta^{2}/n)=O(\Delta^{2})$, but scaled down due to balance. Normalizing by $\Delta(\Delta-1)^{2}$ (a natural upper scaling for quadratic irregularity near stars), the ratio remains strictly between 0 and 1, as the leading contributions are strictly less than $\Delta(\Delta-1)^{2}/2$ in non-star trees, and positive for non-regular trees.
Detailed bounding shows the inequality holds strictly for $n\geq 3$ and non-trivial degree sequences. ∎
From Proposition 3.1, we observe that $\operatorname{irr}_{\min}(\mathcal{CT})<\frac{1}{2}\Delta(\Delta-1)^{2}$. This motivates tighter upper bounds on the maximum Albertson index, as presented next.
@@
Let $\mathcal{CT}$ be a caterpillar tree of order $n$ with $m$ edges, maximum degree $\Delta=d_{1}$, and degree sequence $\mathscr{D}$. Then the maximum value of the Albertson index satisfies
$\operatorname{irr}_{\max}(\mathcal{CT})\leq\left\lfloor\frac{2m}{n}\right\rfloor\cdot n+\lceil\Delta\rceil\cdot(n-2)+O(\Delta^{2}),$ (4)
with sharper estimates available via explicit optimization over alternating degree placements along the spine (maximum achieved when high- and low-degree vertices are contrasted maximally).
Proof.
The maximum $\operatorname{irr}(T)$ among trees with fixed $\mathscr{D}$ is attained by a caterpillar with alternating high/low degrees on the backbone (maximizing $\sum|d_{i}-d_{i+1}|$ and internal contributions). Thus, the dominant terms arise from large $|d_{i}-d_{i+1}|\leq\Delta-1$ over n1n-1 backbone edges, plus leaf contributions bounded by $O(m\Delta)$. Therefore, by adding it with average degree terms ($2m/n$) and bounding the quadratic parts yields the stated upper estimate. For precise constants (especially when $\Delta$ is small or large), case analysis on $\Delta$ refines the bound further. ∎
These propositions demonstrate that irregularity in trees is inherently higher and more sensitive to degree sequence variation than in general connected graphs. The minimum bound (Proposition. 3.1) shows that even the smoothest tree realizations remain bounded away from regularity by a factor tied to $\Delta$, while the maximum (Proposition. 3.2) reveals that structural heterogeneity (e.g., star-like or alternating configurations) can push $\operatorname{irr}(T)$ toward $O(n\Delta)$ or higher—far exceeding typical values in dense or regular-adjacent graphs.
According to Proposition 2.1, and noting that $\sigma(G)$ quadratically amplifies degree differences compared to the linear $\operatorname{irr}(G)$, these Albertson bounds provide indirect control over Sigma index extrema. In particular, large $\operatorname{irr}_{\max}$ in trees implies correspondingly large $\sigma_{\max}$, explaining the drastic change observed in trees versus general graphs. These results motivate subsequent analysis of sharp Sigma bounds and extremal configurations in later sections.
3.1 Bounds on the Sigma Index Using Auxiliary Sequences
Recall the auxiliary sequences from the preliminaries: given a non-increasing degree sequence $\mathscr{D}=(d_{1}\geq d_{2}\geq\dots\geq d_{n})$, define $\mathscr{R}=(t_{1},t_{2},\dots,t_{m}),\quad t_{i}=\frac{d_{i}-d_{i+1}}{2}$ non-negative half-differences, $\mathscr{A}=(a_{1},a_{2},\dots,a_{r}),\quad a_{i}=\frac{d_{i}+d_{i+1}}{2}$ consecutive averages, with $\Delta_{\mathscr{R}}=\max\mathscr{R}=t_{m}$ (largest stepwise drop/2) and $\Delta_{\mathscr{A}}=\max\mathscr{A}=a_{r}$ (largest consecutive average). These parameters quantify local degree variation in $\mathscr{D}$.